OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? -five thirds five thirds -three fifths three fifths

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 0- 3)/(-3-2) m = (-3)/(-5) m = 3/5 So the slope of the line that passes through the two points is 3/5